# Through the eyes of a reader and science communicator: science in the mainstream and in the genre literature of yesterday and today

### 29 Dec 2023 v2

@@ Abstract
For most writers the science is either an exotic setting or a source of thrilling conflict that would drive the story forward. For a communicator it is the other way around – the science is neatly wrapped in a package of literary tools that make it “invisible” while it remains tangible and most importantly – it can be conveyed to the reader in understandable terms. There are many examples showing how these seemingly contradicting goals can complement each other successfully. I will review how the science was communicated by mainstream and genre writers of yesterday and today, and in different (not necessarily anglophone) cultures. I will bring forward the best and the worst examples that illuminate various astronomical concepts. Finally, I will discuss how we can use them both in outreach and education. Contrary to many similar summaries I will concentrate on some often overlooked mainstream literary examples, including the plays The Physisists by Friedrich Dürrenmatt and Copenhagen by Michael Frayn, the novel White Garments by Vl. Dudintsev and even an episode of the Inspector Morse TV show, featuring scientists. I will also mention in passing a few less well known genre books.
@@ 1 Introduction
2 Scientists though the eye of the artists
Scientists are often featured in popular arts. To remind a few, almost household names: Susan Calvin, Indiana Johns, Victor Frankenstein, Dr. Strangelove, Capt. Nemo, Emmett Brown, Eleanor Arroway, Spock, Dr. No, Sheldon Cooper, Dana Scully, Dr. Kelvin, and even Sherlock Holmes. I have intentionally included in the the list both males and females, protagonists and antagonists, to demonstrate the diversity of the characters, the variety of their motives and the range of their actions. A lesser known name is that of Dr. Kelvin, the psychologist from the novel Solaris by Stanislaw Lem (and two motion pictures based on this book).
@@ 2 Scientists though the eye of the artists
For writers and screenwriters invoking scientists is often an easy way to create all-powerful menacing figures that would drive the story forward without the need to invest much though and imagination. Further fuel into this negative outlook to science is provided by two notable novels addressing the devastating consequences that the lack of scientific integrity can bring – Solar by Ian McEwan and The White Garments by Vladimir Dudintsev. The former has a modern setting and follows the life story of an eco-scientist who makes some very questionable decisions. While at its core this is a novel about integrity, the topic is shown to exceeds the limits of the guarded and somewhat confined scientific environment, it is shown as a problem of world shattering proportions. The second book appeared in the first years of the Soviet Perestroyka, when Eastern Europe opened up and faced many issues, including to its own past. It describes a conflict dating from the early post-WWII years while Stalin was still in power and the science succumbing to or rather was trying to fight a loosing battle against ideology. The novel asked the uncomfortable question what can one do, if the research suggest that the world pretty much disagrees with the established ideology and with the party line. Silence and internal immigration is one answer. Summarizing, these two books offer less than favorable images of scientists, and they are a bit too big of a reading to be useful topic starters that would introduce students to ecology and biology,
Another example of a fictionalized presentation of scientists in a completely different environment comes from the Bulgarian writer Dimitar Peev (1919-1996) who published in 1966 a space faring novel The Photon Starship. The action is set in a communist Utopia and follows a sub-light speed trip to the nearest star system Alpha Cen. Surprisingly for a books of that era it contains very little ideology and a lot of scientific speculations, some – quite relevant (for how these can be used as educational topic raisers see [3]). The entire expedition consists of highly motivated scientists, so many that nobody really stands out. This was a concession to the politics back them – the individual did not matter, the collective and the society ruled. However, the increasing average number of authors per paper today may be returning us back to this idea. We can safely say that Peev has adopted a science-optimistic stand. His book offers insights into planet formation and stellar binary formation, although critical comparison with the modern understanding of these topics is needed.
@@
However, the novel should be taken with a big grain of salt as a guide to real-life university life, because the author has made some big concessions to build suspense and boost conflicts. Without revealing too much content, I can only tell that Laura, the graduate student protagonist, suffers from stage fright and have difficulties giving presentations. Every institution that I have studies or worked at in my career had a student seminar, where the students would present their work to each other, in a friendly environment, just to overcome this particular problem. Still, Stars and Waves can be a good vehicle to introduce the audience to academic life, provided such points are addressed and discussed. Unlike the books mentioned earlier it offers much more balanced view. The novel is an excellent introduction into a discussion of a wide range of astronomy and astrobiology topics – from astroclimate and telescope design to the searches for bio- and technosignatures.
@@
The most popular TV series has been less than merciful to the scientists they have portrayed. Inspector Morse and it sequel Lewis turn Oxford into a dangerous place with a high crime rate. The academics are frequently involved, often shown in negative light, being involved in gambling, extramarital affairs and now and then – murders. The protagonists often make a point of the hubris that the scientists tend to exhibit both in their work and personal lives (even though detective himself is not completely alien to this character trait). The Big Bang Theory makes fun of the scientist – without having to apply much imagination to the script, because the real life suffices to make the audience laugh. At best the show evokes sympathy to the “poor” scientists who are portrayed unable to coop with the challenges of everyday life. Both these TV shows paint negative portraits of scientists, albeit from very different angles. Inspector Morse can hardly help as a science intro, but The Big Bang Theory often touches various science questions in detail.
@@
The theatrical stage has brought is two notable examples of science portrayals – the plays Copenhagen and The Physicists. The first, written by Michael Frayn, is based on a real-life visit that Werner Heisenberg paid to Niels Bohr and his wife Margrethe Bohr. There are differences in the accounts the participants gave afterwords. Clearly, they discussed the atomic bomb, the implications and the ethics of building it. After the meeting Heisenberg went on to work on the Nazi nuclear program, together with Otto Hahn, Carl Friedrich von Weizsäcker and other prominent physicists. Bohr escaped to UK and later to US. Overall, this is a thoughtful piece that puts forward a picture of the scientific word that is aware of the consequences that the researchers will have to face in their work [^1] The movie Oppenheimer that premiered after the EAS 2023 addresses similar issues and it probably helped the public to understand and sympathize with the scientists more than any of the works mentioned here.. While Copenhagen does touch upon the basic nuclear physics, it is of more help for discussions on scientific ethics.
@@ 3 Discussion and conclusions
Our survey how research scientist are drawn in arts is broad, coarse and certainly not pretending for completeness in any meaningful way. Unfortunately, even this quick overview drives us towards a rather grim conclusion: with a few rare exceptions the researchers are seen either as logical but cold and hubristic beings devoid of human qualities, who can not be relied upon, or as ill-adjusted rueful loners. The power of fiction is to influence people, either on conscious or on subconscious level, and here it works toward undermining the public trust in scholarly figures. The exceptions almost exclusively come from the pen of a scholarly figures and reach limited audience.
@@ Acknowledgements
This is an extended write up of a contributed talk presented at the European Astronomical Society (EAS), Special Session 39 “Sci-Art: Communicating Science Through Art”, held held in Kraków, Poland on Jul 10-14, 2023.

### 28 Apr 2024 v3

@@ Abstract
For most writers the science is either an exotic setting or a source of thrilling conflict that would drive the story forward. For a communicator it is the other way around – the science is neatly wrapped in a package of literary tools that make it “invisible” while it remains tangible and most importantly – it can be conveyed to the reader in understandable terms. There are many examples showing how these seemingly contradicting goals can complement each other successfully. I will review how the science was communicated by mainstream and genre writers of yesterday and today, and in different (not necessarily anglophone) cultures. I will bring forward the best and the worst examples that illuminate various astronomical concepts. Finally, I will discuss how we can use them both in outreach and education. Contrary to many similar summaries I will concentrate on some often overlooked mainstream literary examples, including the plays The Physicists by Friedrich Dürrenmatt and Copenhagen by Michael Frayn, the novel White Garments by Vl. Dudintsev and even an episode of the Inspector Morse TV show, featuring scientists. I will also mention in passing a few less well known genre books.
@@ 1 Introduction
2 Scientists through the eye of the artists
Scientists are often featured in popular arts. To remind a few, almost household names: Susan Calvin, Indiana Jones, Victor Frankenstein, Dr. Strangelove, Capt. Nemo, Emmett Brown, Eleanor Arroway, Spock, Dr. No, Sheldon Cooper, Dana Scully, Dr. Kelvin, and even Sherlock Holmes. I have intentionally included in the list both males and females, protagonists and antagonists, to demonstrate the diversity of the characters, the variety of their motives and the range of their actions. A lesser known name is that of Dr. Kelvin, the psychologist from the novel Solaris by Stanislaw Lem (and two motion pictures based on this book).
@@ 2 Scientists through the eye of the artists
For writers and screenwriters invoking scientists is often an easy way to create all-powerful menacing figures that would drive the story forward without the need to invest much though and imagination. Further fuel into this negative outlook to science is provided by two notable novels addressing the devastating consequences that the lack of scientific integrity can bring – Solar by Ian McEwan and The White Garments by Vladimir Dudintsev. The former has a modern setting and follows the life story of an eco-scientist who makes some very questionable decisions. While at its core this is a novel about integrity, the topic is shown to exceeds the limits of the guarded and somewhat confined scientific environment, it is shown as a problem of world shattering proportions. The second book appeared in the first years of the Soviet Perestroyka, when Eastern Europe opened up and faced many issues, including to its own past. It describes a conflict dating from the early post-WWII years while Stalin was still in power and the science succumbing to or rather was trying to fight a loosing battle against ideology. The novel asked the uncomfortable question what can one do, if the research suggest that the world pretty much disagrees with the established ideology and with the party line. Silence and “internal immigration” [^1] A term used in the former Eastern Europe to signify a person’s refusal to participate in public life; a refusal to use social media is a modern day analog of this behaviour. is one answer. Summarizing, these two books offer less than favorable images of scientists, and they are a bit too big of a reading to be useful topic starters that would introduce students to ecology and biology.
Another example of a fictionalized presentation of scientists in a completely different environment comes from the Bulgarian writer Dimitar Peev (1919-1996) who published in 1966 a space faring novel The Photon Starship. The action is set in a communist Utopia and follows a sub-light speed trip to the nearest star system Alpha Cen. Surprisingly for a book of that era it contains very little ideology and a lot of scientific speculations, some – quite relevant (for how these can be used as educational topic raisers see [3]). The entire expedition consists of highly motivated scientists, so many that nobody really stands out. This was a concession to the politics back them – the individual did not matter, the collective and the society ruled. However, the increasing average number of authors per paper today may be returning us back to this idea. We can safely say that Peev has adopted a science-optimistic stand. His book offers insights into planet formation and stellar binary formation, although critical comparison with the modern understanding of these topics is needed.
@@
However, the novel should be taken with a big grain of salt as a guide to real-life university life, because the author has made some big concessions to build suspense and boost conflicts. Without revealing too much content, I can only tell that Laura, the graduate student protagonist, suffers from stage fright and have difficulties giving presentations. Every institution that I have studied or worked at in my career had a student seminar, where the students would present their work to each other, in a friendly environment, just to overcome this particular problem. Still, Stars and Waves can be a good vehicle to introduce the audience to academic life, provided such points are addressed and discussed. Unlike the books mentioned earlier it offers a much more balanced view. The novel is an excellent introduction into a discussion of a wide range of astronomy and astrobiology topics – from astroclimate and telescope design to the searches for bio- and technosignatures.
@@
The most popular TV series has been less than merciful to the scientists they have portrayed. Inspector Morse and it sequel Lewis turn Oxford into a dangerous place with a high crime rate. The academics are frequently involved, often shown in negative light, being involved in gambling, extramarital affairs and now and then – murders. The protagonists often make a point of the hubris that the scientists tend to exhibit both in their work and personal lives (even though the detective himself is not completely alien to this character trait). The Big Bang Theory makes fun of the scientist – without having to apply much imagination to the script, because the real life suffices to make the audience laugh. At best the show evokes sympathy to the “poor” scientists who are portrayed unable to coop with the challenges of everyday life. Both these TV shows paint negative portraits of scientists, albeit from very different angles. Inspector Morse can hardly help as a science intro, but The Big Bang Theory often touches various science questions in detail.
@@
The theatrical stage has brought us two notable examples of science portrayals – the plays Copenhagen and The Physicists. The first, written by Michael Frayn, is based on a real-life visit that Werner Heisenberg paid to Niels Bohr and his wife Margrethe Bohr. There are differences in the accounts the participants gave afterwords. Clearly, they discussed the atomic bomb, the implications and the ethics of building it. After the meeting Heisenberg went on to work on the Nazi nuclear program, together with Otto Hahn, Carl Friedrich von Weizsäcker and other prominent physicists. Bohr escaped to the UK and later to the US. Overall, this is a thoughtful piece that puts forward a picture of the scientific word that is aware of the consequences that the researchers will have to face in their work [^2] The movie Oppenheimer that premiered after the EAS 2023 addresses similar issues and it probably helped the public to understand and sympathize with the scientists more than any of the works mentioned here.. While Copenhagen does touch upon the basic nuclear physics, it is of more help for discussions on scientific ethics.
@@ 3 Discussion and conclusions
This survey how research scientist are drawn in arts is broad, it lacks resolution and certainly does not claim completeness in any meaningful way. Unfortunately, even this quick overview drives us towards a rather grim conclusion: with a few rare exceptions the researchers are seen either as logical but cold and hubristic beings devoid of human qualities, who can not be relied upon, or as ill-adjusted rueful loners. The power of fiction is to influence people, either on conscious or on subconscious level, and here it works toward undermining the public trust in scholarly figures. The exceptions almost exclusively come from the pen of a scholarly figures and reach limited audience.
@@ Acknowledgements
This is an extended write up of a contributed talk presented at the European Astronomical Society (EAS), Special Session 39 “Sci-Art: Communicating Science Through Art”, held in Kraków, Poland on July 10-14, 2023.